top of page

Bear Creek Fire Update 08/19/2020


August 19, 2020   8:00 a.m. 

CURRENT SIZE:  7,543 acres (acres mapped by overnight infrared flight), with a 10% containment.

PERSONNEL AND RESOURCES ON THE FIRE:  Resources assigned to the fire include eight helicopters, eight Type 2 IA crews, three Type 1 crews, 21 engines, and 480 total personnel.


YESTERDAY’S ACTIVITY:  Fire behavior was quiet yesterday, with the most active area on the southeast corner. Residual heat was internal to the constructed lines. Crews on the north flank continued to reinforce and hold the line along their previous burnout operation. To the west, crews worked to connect line from the 515 road to the Continental Divide, continuing with the multi-day project, with support from heavy equipment. On the south flank, crews continued to improve the fuel break to the Continental Divide. Helicopters and scoopers provided aerial support. The fixed-wing “scooper” airplanes were reassigned mid-day to assist with new fires in other areas of the country. Night/swing shift patrolled the southeast side of the fire. Resource advisors from multiple jurisdictions are communicating about operations on the ground and coordinating rehabilitation needs. 

TODAY’S WEATHER:  Thunderstorms are expected in the fire area this afternoon with a 55% chance of lightning and a 50% chance of wetting rain. Wetting rain could dampen fire activity; however, with thunderstorms, shifting winds from the west/northwest could reach gusts up to 30 mph. Winds over the fire area are highly variable due to topographic features and the proximity to the Continental Divide. Wind could cause spotting in the northeast area of the fire, potentially causing fire runs. The northwest side of the fire could advance further north and will likely see interior fire activity given winds associated with thunderstorms. Fire activity is expected toward to further back down into Black Canyon. Temperatures are expected to be in the 80s with a relative humidity in the low 20s. A dry and windy pattern of weather is expected to return starting tomorrow. Winds and an inversion at helibase could affect aerial support later in the day.  

TODAY’S PLANNED ACTIVITY: Crews will work to improve the line on the 1883 road on the southern flank. To the north, crews will hold and secure line along the 6869 road. On the west side of the fire, crews will improve the fuel break to the Continental Divide. Crews will continue to work to connect line from the 515 road to the Continental Divide with support from heavy equipment. Air support will be adding a secondary heli-well near the mouth of Black Canyon. There will be no night patrol tonight, as those resources have transitioned into the day shift to assist with operations on the active southern flank. 

CLOSURES:  Area, road, and trail closures are in effect. Refer to the “Closures” and “Photos” tabs of the Inciweb site for narrative descriptions and maps of all agency closures. Lemhi Pass Road is closed from its junction with Montana Highway 324 to Lemhi Pass. The public is strongly encouraged to refrain from entering the area/nearby road system due to heavy fire traffic.  


309 views0 comments


bottom of page